OpenStudy (anonymous):

If the price of tea is increased by 1% how much percent must a housewife reduce her consumption of tea, to have no extra expenditure?

OpenStudy (anonymous):

\[\frac {1}{100+1}\times 100 =0.990099 \approx 0.1\%\]

OpenStudy (anonymous):

pot+(pot*.01)= npot pot is price of tea npot is new price

OpenStudy (anonymous):

huh????

OpenStudy (anonymous):

npot-(pot*.01)= amnt she should reduce here tea consumption by based on price.

OpenStudy (anonymous):

i think in percent form it be 100* (pot/npot)